OpenStudy (anonymous):

Estimate by rounding to the nearest half dollar. $10.13 + $1.46 = 1) $10.00 + $2.00 = $12.00 2) $10.00 + $1.50 = $11.50

alones (alones):

Just simplify 10.13+1.46 =10.13+1.46 =11.59

OpenStudy (anonymous):

Thanks :)

alones (alones):

mhm yw hehe ;]